OpenStudy (anonymous):

Find the y-intercept of the line 3x-4y=12

OpenStudy (anonymous):

Shut up.

OpenStudy (anonymous):

X=0 y = -3 (0,-3)

OpenStudy (anonymous):

don't help him if he's not going to try to learn how to do them. we aren't a cheat resource, we are here to help people learn.

OpenStudy (anonymous):

3x-4y=12 -4y=12-3x y=-3+(3/4)x or y=(3/4)x+-3 So, y-intercept = -3

OpenStudy (anonymous):

Find the y-intercept of the line 3x-4y=-12

OpenStudy (anonymous):

-3